OpenStudy (anonymous):

Write the complex number in trigonometric form. z=3i

OpenStudy (unklerhaukus):

\[z=3i\]\[=3(\cos(\theta)+i\sin(\theta))\]\[=3(\cos(π/2)+i\sin(π/2)) \]\[=3i\sin(π/2)\]\[=3\text{cis}(π/2)\]

OpenStudy (unklerhaukus):

oh, and keep batmaning ,